You are making a $120,000 investment and feel that a 14% rate of return is reasonable, given the nature of the risks involved. You expect to receive $48,000 in the first year, $54,000 in the second year, and $76,000 in the third year. You expect! to pay out $12,000 as a disposal cost in the fourth year. What is the net present value of this investment given your expectations? 
Answer #2

NPV = -$120,000 + $48,000/1.14 + $54,000/1.142 + $76,000/1.143 + (-$12,000/1.144)

NPV = $7,849.38
Answer #1

Net present value is equal to present value of cash inflows- present value of cash outflows

= -120,000 + 48000*PVF(14%, 1 year) +54000*PVF(14%, 2 Years) + 76,000*PVF(14%, 3 years) - 12,000*PVF(14%, 4 years)

=-120,000 +48,000*0.877 + 54,000*0.769 + 76000*0.675 -12000*0.592

=-120,000+42096+41526+51300 - 7104

=$7,818